# CoRA: Collaborative Information Perception by Large Language Model’s Weights for Recommendation

### v2

@@ Abstract
Involving collaborative information in Large Language Models (LLMs) is a promising technique for adapting LLMs for recommendation. Existing methods achieve this by concatenating collaborative features with text tokens into a unified sequence input and then fine-tuning to align these features with LLM’s input space. Although effective, in this work, we identify two limitations when adapting LLMs to recommendation tasks, which hinder the integration of general knowledge and collaborative information, resulting in sub-optimal recommendation performance. (1) Fine-tuning LLM with recommendation data can undermine its inherent world knowledge and fundamental competencies, which are crucial for interpreting and inferring recommendation text. (2) Incorporating collaborative features into textual prompts disrupts the semantics of the original prompts, preventing LLM from generating appropriate outputs. In this paper, we propose a new paradigm, CoRA (an acronym for Collaborative LoRA), with a collaborative weights generator. Rather than input space alignment, this method aligns collaborative information with LLM’s parameter space, representing them as incremental weights to update LLM’s output. This way, LLM perceives collaborative information without altering its general knowledge and text inference capabilities. Specifically, we employ a collaborative filtering model to extract user and item embeddings, converting them into collaborative weights with low-rank properties through the collaborative weights generator. We then merge the collaborative weights into LLM’s weights, enabling LLM to perceive the collaborative signals and generate personalized recommendations without fine-tuning or extra collaborative tokens in prompts. Extensive experiments confirm that CoRA effectively integrates collaborative information into LLM, enhancing recommendation performance.
@@ Introduction
(2) Existing methods endeavor to align collaborative information with the input space of LLMs by embedding collaborative features of users and items into textual prompts. Nevertheless, this practice can disrupt the LLM’s comprehension of the original text semantics. For instance, as shown in Fig. 2, while we expect Vicuna-7B without fine-tuning to repeat user and item description, the revised prompts fail to guide the LLM to generate the correct responses. Overall, existing studies limit general knowledge and collaborative information integration when adapting LLM to recommendation tasks.
@@
Specifically, the prediction for each user-item pair is divided into two parts: collaborative weights generation and text inference with collaborative weights. First, we utilize a pre-trained collaborative filtering model to extract user and item embeddings, which are then processed by a collaborative weight generator. The generator converts these embeddings to collaborative perceptual weights similar to “low-rank adaption” (LoRA (Hu et al. 2022)). We incorporate the collaborative weights into the LLM’s pre-trained weights to endow it with the capability to perceive collaborative information between the user and the item, allowing it to generate personalized predictions for each user-item pair with general text prompts. Importantly, our technique does not alter the LLM’s understanding of the original text, as it does not introduce new tokens. Instead, it updates LLM’s output with incremental weights. For tasks that do not involve collaborative information, our method seamlessly switches to utilizing the frozen backbone to infer textual prompts. Thereby, our procedure inherently addresses the aforementioned problems. To sum up, our contributions are summarized as follows:
  • We explore a novel paradigm for equipping LLMs with collaborative perception capability by merging collaborative weights into LLM’s pre-trained weights. It can be seen as Collaborative LoRA (CoRA), facilitating the cooperation of text and collaborative information. To the best of our knowledge, we are the first to endeavor collaborative perception in the parameter space of LLMs.
  • Following the CoRA method, we propose a collaborative weights generator that transforms collaborative features extracted from conventional recommenders into collaborative weights through a decoder-only architecture with a low-rank property.
@@
A new trend is surfacing to include collaborative information in LLMs. Some studies focus on discovering user and item encoding methods to introduce new tokens through vocabulary expansion(Zheng et al. 2023; Hua et al. 2023; Rajput et al. 2023; Zheng et al. 2024). Others explore extracting collaborative information using latent factor models and aligning it with the input space of LLMs (Zhang et al. 2023, 2024b; Liao et al. 2023; Zhang and wei Wang 2023; Li et al. 2023d). While these methods exhibit strong performance, they constrain specific general capabilities of LLMs. Moreover, due to the hybrid prompting methods, collaborative features can disrupt original textual semantics. In this study, we propose CoRA, which generates low-rank collaborative weights like LoRA and injects them into LLM’s weights without extra inference overhead.
@@
Figure 3: Details of the LLM Decoder Block. (a) depicts the details of the LLM decoder block. (b) illustrates the multi-head self-attention module.
LLMs refer to a class of language models equipped with billions of parameters. As depicted in Fig. 3, their decoder block contains a multi-head self-attention module and a feed-forward network.
Self-attention contains four types of linear layers: query $W_{Q}$, key $W_{K}$, value $W_{V}$, and output $W_{O}\in\mathbb{R}^{h\times d}$, where $h$ and $d$ are dimensions of the hidden states and attention head. For each input sequence with $L$ tokens $X\in\mathbb{R}^{L\times h}$, it is multiplied by linear layers $W_{Q}$, $W_{K}$, $W_{V}$, obtaining $Q=XW_{Q}$, $K=XW_{K}$ and $V=XW_{V}$. Then, the self-attention operation is performed along the sequence dimension as follows:
$\mbox{Attention}(Q,K,V)=\mbox{softmax}(\frac{QK^{\top}}{\sqrt{d}})V.$ (2)
Then, the outputs parallel in different heads are concatenated and again projected by linear transformation matrix $W_{O}$, resulting in the final values.
Figure 4: Model architecture overview of our CoRA. The left path extracts user and item embeddings using a frozen collaborative model and converts them to collaborative weights through the generator. The right path fills the text fields in the prompt template, introducing textual descriptions for inference. Finally, the collaborative weights are merged into the LLM’s weights, enabling the LLM to perceive collaborative information without any fine-tuning or extra tokens in textual prompts.
Feed-forward Network is an MLP with two fully connected layers and one non-linear activation function. The formulation is as follows:
$\mbox{FFN}(x)=\sigma(xW_{1})W_{2},$ (3)
where $x$ is the input token, $W_{1}$ and $W_{2}$ are the weights of fully connected layers, and $\sigma$ is an activation function. Overall, there are five types of linear weights in a decoder block, including $W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$ in the self-attention module, and $W_{1}$, $W_{2}$ in the feed-forward module.
@@
In this section, we introduce our collaborative weights generator and CoRA method, which effectively integrates collaborative information into LLMs in the parameter space. The overall framework is shown in Fig. 4. First, we explain how to obtain collaborative weights from the pre-trained user and item embeddings through a collaborative weights generator. Then, we elaborate on how to equip LLM with collaborative perception capability with the help of generated collaborative weights, followed by a description of the prediction and training strategy.
Collaborative Weights Generator
Figure 5: Collaborative Weights Generator. Fig. (a) illustrates the pipeline of our collaborative weights generator. We set $k$ learnable perceptual queries, which interact with collaborative features through cross-attention modules to perceive collaborative information. Fig. (b) demonstrates that our approach is conceptually consistent with LoRA.
For a user-item pair $(u,i)$, we first obtain user and item embeddings $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d_{c}}$ from a pre-trained CF model and concatenate them as $[\mathbf{e}_{u},\mathbf{e}_{i}]$. To convert the features to collaborative weights $\Delta W\in\mathbb{R}^{h\times h}$, we propose a collaborative weights generator inspired by BLIP-2 (Li et al. 2023a).
Specifically, we design a decoder-only architecture to generate collaborative weights $\Delta W$. As shown in Fig. 5(a), the weights generator contains $N$ decoder blocks, each consisting of a self-attention module, a cross-attention module, and a feed-forward network. We set the hidden dimension of the generator to be equal to $d_{c}$. In the generator, $k$ learnable perceptual queries absorb collaborative information from different $k$ representation sub-spaces. In each decoder block, perceptual queries first interact with each other through self-attention modules, then interact with frozen collaborative features through cross-attention modules to perceive collaborative information, and finally go through a feed-forward network. Afterwards, we obtain $k$ collaboration-aware features $q_{c}\in\mathbb{R}^{d_{c}}$, which should be mapped and aligned with LLM’s parameter space. However, since $d_{c}\ll h\cdot h$, it will introduce a large number of parameters to directly map the dimension from $d_{c}$ to $h\cdot h$. Accordingly, a low-rank computation is considered in this process to reduce the computation costs. In particular, we adopt a shared linear layer $W_{share}\in\mathbb{R}^{d_{c}\times h\cdot r}$ to project all features $q_{c}\in\mathbb{R}^{d_{c}}$ to $W_{c}\in\mathbb{R}^{h\cdot r}$ as follows:
$W_{c}=q_{c}W_{share}$ (4)
where $W_{c}\in\mathbb{R}^{h\cdot r}$ is collaborative parameter and $r$ is the rank for collaborative weights.
Then, we reshape collaborative parameters $W_{c}$ as $h\times r$ and aggregate collaborative information from different sub-spaces with a mean-pooling operation. To obtain collaborative weights $\Delta W$, we ascend the dimension from $h\times r$ to $h\times h$ through a linear projection $W_{proj}\in\mathbb{R}^{r\times h}$ as follows:
$\Delta W=W_{c}W_{proj}.$ (5)
It is worth noting that the form of this low-rank calculation is equivalent to LoRA (Hu et al. 2022), where $W_{c}$ and $W_{proj}$ are the same as $\Delta W_{A}$ and $\Delta W_{B}$ in LoRA. Consequently, as illustrated in Fig. 5(b), our collaborative weights generator can be seen as a “Collaborative LoRA” generator.
@@
Existing LLMRec methods project user and item embeddings into the LLM’s input space to integrate collaborative and textual information. These approaches interfere with the LLM’s comprehension of the original textual prompts using its general capabilities, resulting in sub-optimal performance. To address this issue, we propose including collaborative information in LLM’s parameter space. Concisely, we convert collaborative information of input user-item pair $(u,i)$ to collaborative weights by the generator and merge them into LLM’s weights to enable LLM to perceive collaborative information without any fine-tuning or extra tokens in textual prompts. Without loss of generality, we utilize Vicuna-7B (Chiang et al. 2023) as the backbone LLM.
Concretely, for an input sample consisting of a user, an item, and their descriptions, we first adopt a CF model $f_{\psi}(\cdot)$ to extract the collaborative features $[\mathbf{e}_{u},\mathbf{e}_{i}]\in\mathbb{R}^{d_{c}}$. Then, we convert these collaborative features to collaborative weights $\Delta W\in\mathbb{R}^{h\times h}$ using the collaborative weights generator. To ensure that LLM can perceive collaborative information while preserving its general language capabilities, we sum $\Delta W$ with the frozen LLM’s weights directly as follows:
$\hat{W}=W+\Delta W=W+W_{c}W_{proj}$ (6)
where $W\in\mathbb{R}^{h\times h}$ denotes the overall weights of LLM.
By absorbing collaborative weights, the LLM inherently acquires the ability to perceive collaborative information. Additionally, this approach avoids interfering with or burdening the LLM’s text processing during inference. We can generate collaborative weights and incorporate them into any type of weights within LLM’s decoder block.
@@
The type of equipped collaborative weights. As mentioned in Sec. Large Language Model, there are five types of weights in LLM’s decoder block, which are query, key, value, output, and feed-forward. We explore the impact of inserting collaborative weights for different types of LLM weights. As shown in Tab. 4, LLM equipped with collaborative weights for all types except feed-forward of weights performs best. Moreover, we observe that the performance of qkvo and qko is much better than qkv, suggesting that the output weights $W_{O}$ are essential for collaborative perception.

### v3

@@ Abstract
Involving collaborative information in Large Language Models (LLMs) is a promising technique for adapting LLMs for recommendation. Existing methods achieve this by concatenating collaborative features with text tokens into a unified sequence input and then fine-tuning to align these features with LLM’s input space. Although effective, in this work, we identify two limitations when adapting LLMs to recommendation tasks, which hinder the integration of general knowledge and collaborative information, resulting in sub-optimal recommendation performance. (1) Fine-tuning LLM with recommendation data can undermine its inherent world knowledge and fundamental competencies, which are crucial for interpreting and inferring recommendation text. (2) Incorporating collaborative features into textual prompts disrupts the semantics of the original prompts, preventing LLM from generating appropriate outputs. In this paper, we propose a new paradigm, Collaborative LoRA (CoRA), with a collaborative query generator. Rather than input space alignment, this method aligns collaborative information with LLM’s parameter space, representing them as incremental weights to update LLM’s output. This way, LLM perceives collaborative information without altering its general knowledge and text inference capabilities. Specifically, we employ a collaborative filtering model to extract user and item embeddings and inject them into a set number of learnable queries. We then convert collaborative queries into collaborative weights with low-rank properties and merge the collaborative weights into LLM’s weights, enabling LLM to perceive the collaborative signals and generate personalized recommendations without fine-tuning or extra collaborative tokens in prompts. Extensive experiments confirm that CoRA effectively integrates collaborative information into LLM, enhancing recommendation performance.
@@ Introduction
(2) Existing methods endeavor to align collaborative information with the input space of LLMs by embedding collaborative features of users and items into textual prompts. Nevertheless, this practice can disrupt the LLM’s comprehenson of the original text semantics. For instance, as shown in Fig. 2, while we expect Vicuna-7B without fine-tuning to repeat user and item description, the revised prompts fail to guide the LLM to generate the correct responses. Overall, existing studies limit general knowledge and collaborative information integration when adapting LLM to recommendation tasks.
@@
Specifically, the prediction for each user-item pair is divided into two parts: collaborative weights generation and text inference with collaborative weights. First, we utilize a pre-trained collaborative filtering model to extract user and item embeddings, which are processed by a collaborative query generator. The generator receives collaborative information and injects it into learnable query embeddings. Inspired by VLoRA (Ma et al. 2024) in computer vision, we transform collaborative queries into collaborative perceptual weights similar to “low-rank adaption” (LoRA (Hu et al. 2022)). We incorporate the collaborative weights into the LLM’s pre-trained weights to endow it with the capability to perceive collaborative information between the user and the item, allowing it to generate personalized predictions for each user-item pair with general text prompts. Importantly, our technique does not alter the LLM’s understanding of the original text, as it does not introduce new tokens. Instead, it updates LLM’s output with incremental weights. For tasks that do not involve collaborative information, our method seamlessly switches to utilizing the frozen backbone to infer textual prompts. Therefore, our procedure inherently addresses the aforementioned problems. To sum up, our contributions are summarized as follows:
  • We explore the issues present when integrating textual prompts and collaborative features in LLM’s input space. To resolve these matters, we propose to equip LLMs with collaborative perception capability by merging collaborative weights into LLM’s pre-trained weights. It can be seen as Collaborative LoRA (CoRA), facilitating the cooperation of text and collaborative information. To the best of our knowledge, we are the first to endeavor collaborative perception in the parameter space of LLMs.
  • Following the CoRA method, we propose a collaborative query generator that injects collaborative features extracted from conventional recommenders into learnable query embeddings and then transforms them into LLM weights space with a low-rank property.
@@
A new trend is surfacing to include collaborative information in LLMs. Some studies focus on discovering user and item encoding methods to introduce new tokens through vocabulary expansion(Zheng et al. 2023; Hua et al. 2023; Rajput et al. 2023; Zheng et al. 2024). Others explore extracting collaborative information using latent factor models and aligning it with the input space of LLMs (Zhang et al. 2023, 2024b; Liao et al. 2023; Zhang and wei Wang 2023; Li et al. 2023d). While these methods exhibit strong performance, they constrain specific general capabilities of LLMs. Moreover, due to the hybrid prompting methods, collaborative features can disrupt original textual semantics. Motivated by VLoRA (Ma et al. 2024), which enables LLMs to perceive visual features by representing visual information as model weights, we propose CoRA, which generates low-rank collaborative weights like LoRA and injects them into LLM’s weights without extra inference overhead.
@@
Figure 3: (a) Architecture of the LLM’s Decoder Block. (b) Details of the multi-head self-attention module.
LLMs refer to a class of language models equipped with billions of parameters. Due to its superiority in performance and training efficiency in generative tasks, LLMs with decoder-only architecture have become mainstream. As depicted in Fig. 3, their decoder block contains a multi-head self-attention module, two add-and-norm modules, and a feed-forward network (Vaswani et al. 2017).
Multi-Head Self-Attention module consists of four types of weights: $W^{Q}$, $W^{K}$, $W^{V}$, and $W^{O}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{head}}}$, where $d_{\text{model}}$ and $d_{\text{head}}$ are dimensions of the input embeddings and attention heads, and $d_{\text{head}}$=$d_{\text{model}}/N_{\text{head}}$. For an input sequence with $L$ tokens $X\in\mathbb{R}^{L\times d_{\text{model}}}$, the calculation of the multi-head self-attention layer is as follows:
$\mbox{MultiHead}(Q,K,V)$ $=\mbox{Concat}(\mbox{head}_{1},\dots,\mbox{head}_{N_{\text{head}}})W^{O}$
$\mbox{where}\ \mbox{head}_{i}$ $=\mbox{Attention}(Q_{i},K_{i},V_{i})$ (2)
$=\mbox{softmax}(\frac{XW^{Q}_{i}XW^{K^{\top}}_{i}}{\sqrt{d_{\text{head}}}})XW^{V}_{i}.$
Figure 4: Model architecture overview of our CoRA. The left path extracts user and item embeddings using a CF model and generates collaborative queries. The right path fills the text fields in the prompt template, introducing textual descriptions for inference. Finally, the collaborative queries are projected into the LLM’s parameter space and merged into the LLM’s weights, enabling the LLM to perceive collaborative information without any fine-tuning or extra tokens in textual prompts.
Feed-forward Network is an MLP with two linear layers and one non-linear activation function as follows:
$\mbox{FFN}(X)=\sigma(XW_{\text{up}})W_{\text{down}},$ (3)
where $W_{\text{up}}$ and $W_{\text{down}}$ are the weights of linear layers, and $\sigma$ is an activation function. Overall, there are six types of linear weights in a decoder block: $W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$ in the multi-head self-attention module, and $W_{\text{up}}$, $W_{\text{down}}$ in the feed-forward module.
@@
In this section, we introduce our collaborative weights generator and CoRA method, which effectively integrates collaborative information into LLMs in the parameter space. The overall framework is shown in Fig. 4. First, we explain how to obtain collaborative queries from the pre-trained user and item embeddings through a collaborative query generator. Then, we elaborate on how to equip LLM with collaborative perception capability with the help of generated collaborative queries, followed by a description of the prediction and training strategy.
Generating Collaborative Queries
For a user-item pair $(u,i)$, we first obtain user and item embeddings $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d_{c}}$ from a pre-trained CF model and concatenate them as $[\mathbf{e}_{u},\mathbf{e}_{i}]$. Then, we fed them into a collaborative query generator to obtain user-item-specific collaborative queries, which bridge the collaborative information and the LLM’s parameter space. Similar to the Q-Former in BLIP-2 (Li et al. 2023a) and the perceptual weights generator in VLoRA (Ma et al. 2024), we initialize $k$ learnable query embeddings and input them into $N$ decoder blocks with cross-attention modules, to absorb collaborative information from different $k$ semantic sub-spaces.
Specifically, as shown in Fig. 4, $k$ query embeddings first interact with each other in the self-attention module, capturing the underlying relationship between different representation sub-spaces. Then, these $k$ query embeddings perceive and comprehend collaborative information in pre-trained user and item embeddings with the help of the cross-attention module. Finally, the $k$ query embeddings are transformed into $k$ deep collaborative features after passing through the feed-forward network. Afterward, we adopt a pooling operation to aggregate collaborative information from $k$ sub-spaces, obtaining the eventual collaborative information-aware query embedding $\mathbf{q}_{c}\in\mathbb{R}^{2d_{c}}$, where $d_{c}$ is the dimension of pre-trained user/item embeddings. In BLIP-2, the output query embedding is linearly projected into the same dimension as the text embedding of the LLM. However, to avoid the collision between textual prompts and collaborative information, we utilize it to generate an incremental weight to guide LLM in recognizing collaborative information and generating appropriate results.
@@
Existing LLMRec methods project user and item embeddings into the LLM’s input space to integrate collaborative and textual information. These approaches interfere with the LLM’s comprehension of the original textual prompts using its general capabilities, resulting in sub-optimal performance. Consequently, we suggest avoiding the incorporation of user/item embeddings and textual prompts. Inspired by the success of VLoRA (Ma et al. 2024) in computer vision, we address this issue by treating the collaborative information as incremental weights of pre-trained LLMs.
By establishing a bridge between collaborative information and the LLM’s parameter space through query embedding $\mathbf{q}_{c}$, our objective is to generate collaborative weights $W_{c}$ that can be seamlessly integrated into the pre-trained LLM weights. As we mentioned in the preliminaries, the shape of LLM weights should be $d_{\text{model}}\times d_{\text{model}}$. With transforming $\mathbf{q}_{c}\in\mathbb{R}^{2d_{c}}$ into $\Delta W\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ directly, we will introduce a transformation matrix $W\in\mathbb{R}^{2d_{c}\times d_{\text{model}}\cdot d_{\text{model}}}$, whose parameter cost is unacceptable. Accordingly, a low-rank computation is considered in this process to reduce the computation costs. Inspired by LoRA (Hu et al. 2022), we utilize a Fully Connected layer to map the output query embedding $\mathbf{q}_{c}$ into LLM’s parameter space as $\Delta W_{A}\in\mathbb{R}^{d_{\text{model}}\times r}$ in LoRA, and ascend the dimension with a final linear projector $W_{\text{proj}}\in\mathbb{R}^{r\times d_{\text{model}}}$, which is equivalent to $\Delta W_{B}$ in LoRA. Eventually, the user-item-specific collaborative weight is calculated by multiplying $\Delta W_{A}$ and $\Delta W_{B}$, which is formulated as:
$W_{c}$ $=\Delta W_{A}\Delta W_{B}$ (4)
$=\mbox{R}(\mathbf{q}_{c}W_{\text{FC}})W_{\text{proj}},$
where $\mbox{R}(\cdot)$ represent a reshape operator, which reshape the product of $\mathbf{q}_{c}$ and $W_{\text{FC}}$ from $\mathbb{R}^{d_{\text{model}}\cdot r}$ into $\mathbb{R}^{d_{\text{model}}\times r}$.
Generally, for an input sample consisting of a user, an item, and their textual descriptions, we extract their collaborative features through a pre-trained CF model and inject the features into a set number of learnable queries with the cross-attention module. Then, we map the collaborative queries into LLM’s parameter space and transform them into incremental LLM weights with low-rank properties. Finally, the collaborative weights can be directly merged into pre-trained LLM weights:
$\hat{W}=W+W_{c}=W+\mbox{R}(\mathbf{q}_{c}W_{\text{FC}})W_{\text{proj}},$ (5)
where $W\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ denotes the overall weights of LLM. By absorbing collaborative weights, the LLM inherently perceives collaborative information without altering the LLM’s understanding of the original textual prompts.
@@
Hyperparameter Settings. Regarding hyperparameter tuning, we explore the learning rate within the range of $[1e-2,1e-3,1e-4]$ and the weight decay within the range of $[1e-2,\dots,1e-6]$. Finally, we train the collaborative weights generator with a learning rate of $21e-$ 2 on Amazon-Book and $31e-$ 3 on ML-1M. On two datasets, the warm-up learning rate is set to $51e-$ 5. The rank $r$ of the collaborative weights is set to 16, and we insert collaborative weights into the weights of four types of LLMs, namely $[W_{Q},W_{K},W_{V},W_{O}]$ after comparing the performances. For all pre-training collaborative filtering models, we set the embedding size to $256$. The learning rate is explored within the range of $[1e-1,\dots,1e-5]$, and the weight decay is explored within the range of $[1e-2,\dots,1e-6]$.
Besides, we adopt an early-stop strategy if AUC on the validation set no longer increases after $20$ epochs to avoid overfitting and achieve the best performance. Our source code included in the Supplementary material provides detailed hyper-parameter settings.
@@
The type of equipped collaborative weights. As mentioned in preliminaries, there are six types of weights in an LLM’s decoder block, which are query, key, value, output, and up&down (feed-forward). We explore the impact of inserting collaborative weights for different types of LLM weights. As shown in Tab. 4, LLM equipped with collaborative weights for all types except feed-forward of weights performs best. Moreover, we observe that the performance of qkvo and qko is much better than qkv, suggesting that the output weights are essential for collaborative perception.